# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v11

@@ D.8 Alternative Identification Strategy
[…]
@@
Figure D.4: Trends in the Two Mining Areas in Iwate and Akita
@@
Notes: This figure shows the time-series plots of the outcomes of the treatment and control groups in Iwate and Akita prefectures used in the regressions shown in Table D.4. Figures D.4aD.4bD.4cD.4e, and D.4d show the average number of people, average crude marriage rate, average marital fertility rate, average sex ratio and average household size, respectively. The data on the IMR in 1925 are unavailable. Source: Created by the author.
@@
Table D.4: Additional Results: Difference-in-Differencesfor Two Coal Mine Areas in Iwate and Akita Prefectures
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses. Notes: This table shows the results for the coal mines subsample in Iwate and Akita prefectures. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively. Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.4 provides the results. In Panel A, column 1 shows the result for the local population, whereas columns 2–5 show those for the marital fertility rate and mortality rates, respectively. In Panel B, columns 1-3 present the results for the early-life mortality rates, whereas columns 4–5 show the results for the sex ratio and household size, respectively. Overall, the DID results align with the findings from my baseline results. First, the opening of these two coal mines increased the local population by approximately $8$%. Although the estimated magnitude is much smaller than the baseline magnitude listed in Table 2636363Notice that I compare my estimates with those for the results in 1935 or 1938 because the DID estimator captures the average treatment effect in the treated municipalities at the post-treatment period., this is rather consistent with the fact that both Ogawa and Miyoshiyama coal mines are considerably small mines with roughly 100 miners. [^64] In fact, this magnitude is rather similar to those estimated for the smaller coal mines in my expanded regressions (Online Appendix Table D.6). The results for the fertility and mortality variables are close to zero and statistically insignificant. This aligns with the small influence of the small-scale mines (Online Appendix D.11). Moreover, both coal mines used few female miners: only 9 and 44 miners in Ogawa and Miyoshiyama coal mines, respectively. Thus, the regulations were less likely to influence these mines. To summarize, the local population growth in both mining areas was mainly led by the influx of male workers, which is supported by the estimate for the sex ratio and average household size (columns 4–5).
@@ D.9 Alternative Minerals: Heavy Metal Mines
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: This table shows the results for the GSC mines sample. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
@@ D.10 Time-Series Plots of the IMR and FDR
Note: This figure shows the national average infant mortality (IMR; solid line) and fetal death rates (FDR; dashed line) in Japan between 1925 and 1940. The IMR is the number of infant deaths per 1,000 live births. The FDR is defined as the number of fetal deaths per 1,000 births. Sources: Created by the author. Data on the number of infant and fetal deaths are from the Statistics Bureau of the Cabinet (1926–1942).
Figure D.5: The Infant Mortality and Fetal Death Rates
between 1925 and 1940
@@ D.11 Heterogeneity with Respect to Scale of Mines
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: MineDeposit ($\leq$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners less than or equal to the median. MineDeposit ($>$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners more than the median. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
@@
  • (66) Social Welfare Bureau. Shussan, shusshō, shizan oyobi nyuȳōjisibō tōkei (The statistics of birth, live births, fetal deaths, and infant and child deaths). [in Japanese]. Social Welfare Bureau, Tokyo, 1941.

### v12

@@ D.8 Alternative Identification Strategy
$\footnotesize{\begin{split}\textit{y}_{i,t}=\pi+\delta\textit{D}_{i,t}+\mathbf{x}^{\prime}_{i,t}\bm{\phi}+\eta_{i}+\lambda_{t}+\omega_{i,t},\end{split}}$ (21)
where $\eta$, $\lambda$, and $\omega$ are the municipal fixed effect, year fixed effects, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
@@
Notes: This figure shows the time-series plots of the outcomes of the treatment and control groups in Iwate and Akita prefectures used in the regressions shown in Table D.4. Figures D.4aD.4bD.4cD.4e, and D.4d show the average number of people, average crude marriage rate, average marital fertility rate, average sex ratio and average household size, respectively. The data on the IMR in 1925 are unavailable.
Source: Created by the author.
@@
Table D.4: Additional Results: Difference-in-Differences
for Two Coal Mine Areas in Iwate and Akita Prefectures
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses.
Notes: This table shows the results for the coal mines subsample in Iwate and Akita prefectures. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively.
Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.4 provides the results. In Panel A, column 1 shows the result for the local population, whereas columns 2–5 show those for the marital fertility rate and mortality rates, respectively. In Panel B, columns 1-3 present the results for the early-life mortality rates, whereas columns 4–5 show the results for the sex ratio and household size, respectively. Overall, the DID results align with the findings from my baseline results. First, the opening of these two coal mines increased the local population by approximately $8$%. Although the estimated magnitude is much smaller than the baseline magnitude listed in Table 2656565Notice that I compare my estimates with those for the results in 1935 or 1938 because the DID estimator captures the average treatment effect in the treated municipalities at the post-treatment period., this is rather consistent with the fact that both Ogawa and Miyoshiyama coal mines are considerably small mines with roughly 100 miners. [^66] In fact, this magnitude is rather similar to those estimated for the smaller coal mines in my expanded regressions (Online Appendix Table D.6). The estimates for fertility and mortality rates are also smaller than the baseline estimates, albeit the fertility pathway is still observed in the statistical sense. This aligns with the small influence of the small-scale mines (Online Appendix D.11). Moreover, both coal mines used few female miners: only 9 and 44 miners in Ogawa and Miyoshiyama coal mines, respectively. Thus, the regulations were less likely to influence these mines. To summarize, the local population growth in both mining areas was mainly led by the influx of male workers, which is supported by the estimate for the sex ratio and average household size (columns 4–5).
@@ D.9 Alternative Minerals: Heavy Metal Mines
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: This table shows the results for the GSC mines sample. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
@@ D.10 Time-Series Plots of the IMR and FDR
Note: This figure shows the national average infant mortality (IMR; solid line) and fetal death rates (FDR; dashed line) in Japan between 1925 and 1940. The IMR is the number of infant deaths per 1,000 live births. The FDR is defined as the number of fetal deaths per 1,000 births.
Sources: Created by the author. Data on the number of infant and fetal deaths are from the Statistics Bureau of the Cabinet (1926–1942).
@@ D.11 Heterogeneity with Respect to Scale of Mines
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: MineDeposit ($\leq$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners less than or equal to the median. MineDeposit ($>$ median) is an indicator variable for the municipalities within 5km from a mine, which had the number of miners more than the median. Panels A and B show the results for the 1930 and 1935 samples, respectively. Column 1 shows the results for the log-transformed population. Column 2 shows the results for the marital fertility rate. Columns 3 and 4 show the results for the male and female mortality rates, respectively. Columns 5 and 6 show the results for the sex ratio and average household size, respectively.
@@
  • (9)
  • (10) Compilation Committee of the Chronology of Chikuhō Coal Mining Industry. Chikuhōtankō kōgyōshi nenpyō (Chronology of Chikuhō Coal Mininig Industry). [in Japanese]. West Japan Cultural Association, Fukuoka, 1973.
@@
  • (34) Fukuoka Prefecture. Fukuokaken tōkeisho, shōwa gonen daisanhen (Fukuoka Prefectural Statistics, volume 3 of 1930 edition). [in Japanese]. Fukuoka Prefecture, Fukuoka, 1932.
  • (35)
@@
  • (54) Yoshihiro Ogino. Chikuhōtankō rōshi kankeishi (History of Labor-Management Relations in Chikuhō Coal Mine). [in Japanese]. Kyushi University Press, Fukuoka, 1993.
  • (55)
@@
  • (72) Social Welfare Bureau. Shussan, shusshō, shizan oyobi nyūyōjisibō tōkei (The statistics of birth, live births, fetal deaths, and infant and child deaths). [in Japanese]. Social Welfare Bureau, Tokyo, 1941.
@@
  • (76) Naoki Tanaka. Kindainihon tankōrōdōshi kenkyū (Research on the History of Coal Mining Labor in Modern Japan). [in Japanese]. Sōfuk̄an, Tokyo, 1984.
  • (77)